[figure: 9.7 cm]

SPIN EXCITATIONS IN DOPED MANGANITES

Andrzej M. Oleś^{∗} and Louis Felix Feiner^{†}

∗ Institute of Physics, Jagellonian University, Reymonta 4, PL-30059 Kraków, Poland † Institute for Theoretical Physics, Utrecht University, NL-3584 CC Utrecht, and Philips Research Labs, Prof. Holstlaan 4, NL-5656 AA Eindhoven, The Netherlands

We investigate a novel effective t−J model for doped La_{1−x}A_{x}MnO_{3} compounds, where A are divalent alkaline earth ions, and 0 < x < 0.5. In the presence of strong Coulomb interaction U at the Mn ions, the superexchange (J-term) between spins S = 2 in the undoped LaMnO_{3} compound follows from virtual d_{i}^{4}d_{j}^{4} ⇀ ↽ d_{i}^{3}d_{j}^{5} excitations due to the hopping of either e_{g} or t_{2g} electron [1]. The hopping of doped holes in La_{1−x}A_{x}MnO_{3} (t-term) occurs in the correlated e_{g} band and depends on the magnetic order according to the double exchange mechanism. It promotes a transition from the A-type antiferromagnetic to a ferromagnetic phase, and modifies the e_{g}-superexchange. Using Schwinger bosons for the S = 2 spins and slave bosons for the two e_{g} orbitals, d_{x^{2}−y^{2}} and d_{3z^{2}−r^{2}}, we derive the magnetic excitations in the ferromagnetic phase. The momentum dependence of the magnon spectrum is determined by the geometry of the perovskite lattice, giving a new interpretation to the earlier result derived for a nondegenerate band [2]. Both the effective ferromagnetic interactions and the stiffness constant increase rapidly with the actual hole doping level. Taking realistic values of the e_{g}-hopping t and Hund’s exchange J_{H} [1], we estimate the magnon dispersion in ferromagnetic doped manganites.

[1] L. F. Feiner and A. M. Oleś, Phys. Rev. B. 59, 3295 (1999).

[2] N. Furukawa, J. Phys. Soc. Jap. 65, 1174 (1996).

[figure: 13.4 cm]

Subject category :

1. Correlated Electrons and High Temperature Superconductors

Presentation mode : oral

Corresponding author : Andrzej M. Oleś

Address for correspondence : Institute of Physics, Jagellonian University, Reymonta 4, PL-30059 Kraków, Poland

Email address : amoles@jetta.if.uj.edu.pl